A conveyor belt costs $40,000 and lasts for three years. A reinforced conveyer belt costs $50,000 and lasts for four years. Assuming an APR of 10%, which type of belt is less expensive?
Please show all your work.
Answer #1

EUAC of first belt = 40000*(A/P, 10%,3) = 40000 * 0.402115 = 16084.59

EUAC of reinforced belt = 50000*(A/P, 10%,4) = 50000 * 0.315471 = 15773.54

As EUAC of reinforced belt is less, it should be selected